OpenStudy (anonymous):

Company ABC produces widgets. They have found that the cost, c (x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $23 to produce 2 widgets, $55 to produce 4 widgets, and $247 to produce 10 widgets. What is the total cost of producing

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

you know the function looks like \[f(x) = ax^2+bx+c\] \[f(2)= 23= 4a+2b+c\] \[f(4)=55=16a+4b+c\] \[f(10)=247=100a+10b+c\]

OpenStudy (anonymous):

you have three equations in three unknowns to solve

OpenStudy (anonymous):

A. $103 B.$153 C.$167D. $177